OpenStudy (yolo_king):

There are 4 blue pencils, 2 red pencils, and 3 orange pencils. What is the probability of choosing an orange pencil then a blue pencil without replacement? A) 2/36 B) 3/12 C) 10/72 D) 12/72

OpenStudy (yolo_king):

I just need help finding these kinds of answers.

OpenStudy (yolo_king):

I think its \[3/9 *4/6\]

OpenStudy (yolo_king):

12/18 = 6/9 = 2/3

OpenStudy (yolo_king):

thats all i got...

OpenStudy (sweetburger):

Ok so your first number is 3/9. How did you get 4/6 for your second number?

OpenStudy (yolo_king):

four blues and it said without replacment so theres six left

OpenStudy (yolo_king):

since i cant replace the oranges

OpenStudy (sweetburger):

but you only removed one pencil so how did you go from 9 total pencils to 6?

OpenStudy (yolo_king):

ooooooooooo i removed one not three my bad so it should be 4/8

OpenStudy (sweetburger):

yea

OpenStudy (yolo_king):

So D?

OpenStudy (calculusxy):

\(\frac{1}{3} \times \frac{1}{2} = \frac{1}{6}\) which is equivalent to \(\frac{12}{72}\. So you're correct.

OpenStudy (calculusxy):

\(\frac{12}{72}\)

OpenStudy (yolo_king):

I got a couple more if you guys dont mind

OpenStudy (yolo_king):

Thanks a lot btw

undeadknight26 (undeadknight26):

I knew the answer was 12/72 I was just confused on how I got there and after I got there once I couldn't get there again it was depressing.